Jim can paint a house in 10 hours and Bill can paint the same house in 12 hours.How long will it take if they work together. Thanks!!
Answer by checkley77(12844)

1/10+1/12=1/x
(12+10)/10*12=1/x
22/120=1/x
22x=120
x=120/22
x=5.455 hours to paint the house working together.